Why can't I get this? I keep getting two different answers
y = -2/5x + 11
y = 2/5x - 11/10

Write the equation in the slope-intercept form. 4x - 10y = 11
Please help.
Ellen

Found 2 solutions by mangopeeler07, stanbon:
Answer by mangopeeler07(462)

Your second answer is correct:

4x-10y=11

Move x to the other side
-10y=-4x+11

Divide both sides by -10
y=4/10x-11/10

Reduce your fractions
y=2/5x-11/10

Answer by stanbon(75887)

y = -2/5x + 11
y = 2/5x - 11/10
--------
Add to solve for "y":
2y = 121/10
y = 121/20
---------------
Solve for "x"
y = (-2/5)x + 11
(121/20) = (-2/5)x + 11
(-2/5)x = 11/10
-2x = 11/2
x = -11/4
================================
Write the equation in the slope-intercept form. 4x - 10y = 11
10y = 4x-11
y = (2/5)x - (11/10)
==========================
Cheers,
Stan H.